# thanks: corresponding authorthanks: corresponding author

### v3

@@ Abstract
Post-selection strategies that discard low-confidence computational results can significantly improve the effective fidelity of quantum error correction at the cost of reduced acceptance rates, which can be particularly useful for offline resource state generation. Prior work has primarily relied on the “logical gap” metric with the minimum-weight perfect matching decoder, but this approach faces fundamental limitations including computational overhead that scales exponentially with the number of logical qubits and poor generalizability to arbitrary codes beyond surface codes. We develop post-selection strategies based on computationally efficient heuristic confidence metrics that leverage error cluster statistics (specifically, aggregated cluster sizes and log-likelihood ratios) from clustering-based decoders, which are applicable to arbitrary quantum low-density parity check (QLDPC) codes. We validate our method through extensive numerical simulations on surface codes, bivariate bicycle codes, and hypergraph product codes, demonstrating orders of magnitude reductions in logical error rates with moderate abort rates. For instance, applying our strategy to the $[[144,12,12]]$ bivariate bicycle code achieves approximately three orders of magnitude reduction in the logical error rate with an abort rate of only 1% (19%) at a physical error rate of 0.1% (0.3%). Additionally, we integrate our approach with the sliding-window framework for real-time decoding, featuring early mid-circuit abort decisions that eliminate unnecessary overheads. Notably, its performance matches or even surpasses the original strategy for global decoding, while exhibiting favorable scaling in the number of rounds. Our approach provides a practical foundation for efficient post-selection in fault-tolerant quantum computing with QLDPC codes.
@@ 1 Introduction
Post-selection can be particularly useful for offline resource state generation processes such as magic state preparation [2, 3], where aborting and retrying failed attempts does not damage existing encoded quantum information. Additionally, when estimating expectation values of observables (e.g., ground-state energies via the variational quantum eigensolver [4, 5]), discarding low-confidence runs can substantially improve estimator fidelity at the cost of additional sampling in the context of quantum error mitigation [6, 7, 8, 9, 10, 11].
@@
The logical gap (defined as the log-likelihood ratio difference between candidate corrections in distinct logical classes) has emerged as a Bayesian-motivated metric quantifying decoding confidence [15, 16, 17, 18]. Physically, according to the statistical-mechanical mapping of error correction, the logical gap of a surface code in the code-capacity setting is proportional to the zero temperature free energy cost of a domain wall in the mapped random-bond Ising model [19, 20]. Post-selection strategies based on the logical gap proved remarkably effective, enabling about 15-fold reductions in error rates with a relative overhead factor of <2<2 [15], error threshold improvements (up to 50% with full post-selection under code capacity noise) [17], and order-of-magnitude reductions in magic state preparation costs [21, 22]. The logical gap has become the standard approach for confidence-based post-selection in QEC.
Nevertheless, despite these advances, the logical gap method faces fundamental limitations that restrict broader applicability. First, computational overhead scales exponentially with the number $k$ of logical qubits (requiring comparative decoding across all $2^{k}$ logical classes), which makes the method prohibitive for codes or circuits involving multiple logical qubits. Second, the method is not guaranteed to work well for general decoders beyond the optimal degenerate maximum likelihood decoder [23, 24] or the minimum-weight perfect matching (MWPM) decoder [25, 26, 27], as a correction within a fixed logical class may fail to represent the likelihood of the class.
@@ 2.1 Soft outputs from decoding
The minimum requirement of a decoder is to infer whether each logical observable needs to be flipped, given the detector outcomes. Here, detectors indicate specific products of measurement outcomes deterministic in the absence of errors, which are normally constructed from two consecutive measurements of a stabilizer. A detector error model (DEM) [47] specifies the prior probabilities of independent fault locations (referred to as “error mechanisms”) as well as the detectors and logical observables flipped by each of them. Decoders primarily aim to address the classification problem: given a DEM and detector outcomes, which logical class the current state belongs to (i.e., which logical observables have been flipped due to errors). In practice, however, decoders typically produce additional information beyond this basic classification. These additional outputs are referred to as the soft outputs of the decoder.
@@
Here we focus on a specific application of soft outputs: quantifying decoding confidence. Specifically, we aim to estimate the reliability of the logical flip inferred by a decoder through analysis of its soft outputs. Given such a confidence metric, we construct a post-selection strategy that accepts a trial only when the metric value exceeds a specified cutoff threshold. We note that soft outputs have found broader applications beyond post-selection, including hierarchical code architectures where confidence information from inner codes guides the decoding of outer codes [16, 29, 53]. However, this work focuses specifically on their application to post-selection.
@@ 2.1.1 Logical gap for quantifying decoding confidence
A prominent approach in the literature for quantifying decoding confidence involves comparative decoding and evaluation of the logical gap (also known as the complementary gap) [15, 16, 17]. Here, comparative decoding refers to a technique that performs decoding within each logical class and selects the minimum-weight correction among all classes. Decoding within a fixed logical class (identified by a vector $\overrightarrow{\mathbf{\lambda}}\in\mathbb{Z}_{2}^{k}$ for the number $k$ of observables) can be implemented by adding $k$ new “virtual” detectors (corresponding to the $k$ observables) into the DEM and assigning specific values (elements of $\overrightarrow{\mathbf{\lambda}}$) to them for decoding. The logical gap $\Delta$ is then defined as the difference between the two smallest log-likelihood weights among the corrections $\{\tilde{E}_{\overrightarrow{\mathbf{\lambda}}}\}_{\overrightarrow{\mathbf{\lambda}}\in\mathbb{Z}_{2}^{k}}$ in different logical classes, i.e.,
$\Delta\coloneqq w_{(2)}-w_{(1)},$
@@
where $w_{(1)}\leq w_{(2)}\leq\cdots\leq w_{(2^{k})}$ are the sorted values of $w(\tilde{E}_{\overrightarrow{\mathbf{\lambda}}})]_{\overrightarrow{\mathbf{\lambda}}\in\mathbb{Z}_{2}^{k}}$. A large logical gap indicates that error patterns in alternative logical classes are significantly less likely than the predicted one, thereby implying high confidence in the decoding outcome.
Despite this general description, the logical gap method has been investigated almost exclusively for the MWPM decoder. This limitation arises from several fundamental problems that prevent broader applicability. First, unless the MWPM decoder is used, we cannot guarantee that decoding within a fixed logical class yields the minimum-weight correction in that class. This means that the weight may not adequately represent the likelihood of the logical class, which could diminish the effectiveness of the logical gap in capturing decoding confidence. Moreover, some decoders may exhibit significantly degraded performance when the logical class is fixed. For example, the union-find (UF) decoder [30] may not be well-suited for comparative decoding, as a cluster that touches a virtual detector (which connects to $d$ or more errors) can grow abnormally fast, thereby failing to accurately capture the underlying clustering structure of the error pattern. Similarly, decoders for QLDPC codes that rely on the low-density property (such as the generalized UF decoder [31]) may also perform poorly due to high-degree virtual detectors with $O(d)$ connections.
@@
While the MWPM decoder does not suffer from these issues, it is applicable only to a limited class of QEC codes, such as repetition and surface codes, where each elementary error mechanism affects at most two detectors [54]. (DEMs may contain error mechanisms affecting more than two detectors, but they can be decomposed into elementary error mechanisms.) A notable exception is the class of two-dimensional color codes [55], which themselves are not compatible with MWPM but allow suboptimal decoders that operate by applying MWPM to specific matchable subgraphs [56, 57, 58, 59, 60, 61, 62]. This makes the logical gap method reasonably effective [22], particularly with the concatenated MWPM decoder [62].
@@
Another significant limitation of the logical gap method is its computational cost, which scales exponentially with the number $k$ of observables, requiring decoding across all $2^{k}$ logical classes. While this scaling may be manageable for memory experiments involving a single logical qubit, it renders the method essentially impractical for scenarios involving multiple logical qubits, e.g., when using codes that encode multiple logical qubits per block or when multiple logical qubits are subject to correlated noise due to multi-qubit gates.
@@ 2.2.1 Clustering-based decoders
Figure 1: Overview of our heuristic cluster-based confidence metrics and post-selection strategies based on them. (a) Given detector outcomes, a clustering-based decoder constructs valid clusters and performs decoding within each cluster in parallel. As an example, we show a distance-9 surface code patch under bit-flip noise. Blue squares denote violated detectors (check nodes), red circles indicate fault nodes within clusters grown by the decoder, and circles marked with $\times$’ represent the final corrections obtained by decoding each cluster. Note that the depicted cluster configuration is provided solely for illustration; actual results may vary depending on the decoder. (b) Cluster statistics are then used to quantify decoding confidence. As the size and number of clusters grow, a larger portion of the logical operator’s support tends to overlap with clusters, thereby increasing the probability of logical failure. Based on this intuition, we define two metrics (the cluster size and LLR norm fractions) in Definitions 1 and 2, and introduce post-selection strategies that abort a trial when the metric value exceeds a chosen cutoff. (c) This approach is further extended to real-time decoding with the sliding-window framework. After decoding each window of $W$ rounds and committing its first $F$ rounds, clusters are constructed from committed errors within the latest LFLF rounds (where $L$ denotes the lookback window size), and a metric is evaluated based on them. If the value exceeds the cutoff, the trial is immediately aborted, thereby avoiding unnecessary computation.
As illustrated in Fig. 1(a), a clustering-based decoder first identifies valid clusters according to a specific rule, which together can explain all violated detectors. These clusters are then decoded individually to produce local corrections, which are then combined into the global correction.
@@
To define terms more rigorously, the check matrix is a binary matrix indicating how error mechanisms (columns) and detectors (rows) are connected, which forms the biadjacency matrix of the bipartite Tanner graph with check nodes (corresponding to detectors) and fault nodes (corresponding to error mechanisms). The fault graph is constructed by projecting the Tanner graph onto fault nodes; namely, a pair of its vertices is connected within the fault graph if the corresponding error mechanisms are commonly involved in at least one detector. A cluster is a set of error mechanisms that forms a connected region in the fault graph. It is considered valid if it contains at least one solution consistent with all detectors that nontrivially involve any error mechanisms in the cluster.
The union-find (UF) decoder is a representative example of this approach applicable to topological codes [30], which works by growing and fusing clusters incrementally until becoming valid (by checking parity), followed by identifying a spanning tree for each cluster (the peeling decoder). The UF decoder was later generalized to quantum low-density parity-check (QLDPC) codes [31]. Another promising strategy involves forming clusters guided by BP outcomes (such that high-likelihood error mechanisms from BP are preferentially included in clusters), which underlies decoders such as BP plus localized statistics decoding (BP+LSD) [32] and ambiguity clustering (AC) [33].
Specifically, the BP+LSD decoder [32] first runs BP, accepting its solution if it converges, and otherwise applying LSD postprocessing. In LSD, clusters are grown incrementally (prioritizing high-likelihood error mechanisms) until they become valid. Each cluster is then decoded by inverting the submatrix of the check matrix restricted to the relevant detectors and a set of linearly independent columns with highest BP likelihoods within the cluster. Crucially, a matrix factorization algorithm called on-the-fly elimination enables efficient validity check and local decoding; since this can be done in parallel on multiple small clusters, BP+LSD can achieve reduced runtime compared to the BP plus ordered statistics decoding (BP+OSD) decoder [51, 49]

### v4

@@ Abstract
Post-selection strategies that discard low-confidence computational results can significantly improve the effective fidelity of quantum error correction at the cost of reduced acceptance rates, which can be particularly useful for offline resource state generation and other moderate-depth fault-tolerant circuits. Prior work has primarily relied on the “logical gap” metric with the minimum-weight perfect matching decoder, but this approach faces fundamental limitations including computational overhead that scales exponentially with the number of logical qubits and poor generalizability to arbitrary codes beyond surface codes. We develop post-selection strategies based on computationally efficient heuristic confidence metrics that leverage error cluster statistics (specifically, aggregated cluster sizes and log-likelihood ratios) from clustering-based decoders, which are applicable to arbitrary quantum low-density parity check (QLDPC) codes. We validate our method through extensive numerical simulations on surface codes, bivariate bicycle codes, and hypergraph product codes, demonstrating orders of magnitude reductions in logical error rates with moderate abort rates. For instance, applying our strategy to the $[[144,12,12]]$ bivariate bicycle code achieves approximately three orders of magnitude reduction in the logical error rate with an abort rate of only 1% (19%) at a physical error rate of 0.1% (0.3%). Additionally, we integrate our approach with the sliding-window framework for real-time decoding, featuring early mid-circuit abort decisions that eliminate unnecessary overheads. Notably, its performance matches or even surpasses the original strategy for global decoding, while exhibiting favorable scaling in the number of rounds. Our approach provides a practical foundation for efficient post-selection in fault-tolerant quantum computing with QLDPC codes.
@@ 1 Introduction
Post-selection can be particularly useful for offline resource state generation processes such as magic state preparation [13, 12], where aborting and retrying failed attempts does not damage existing encoded quantum information. Additionally, when estimating expectation values of observables (e.g., ground-state energies via the variational quantum eigensolver [54, 6]), discarding low-confidence runs can substantially improve estimator fidelity at the cost of additional sampling in the context of quantum error mitigation [9, 58, 15, 68, 47, 14, 71, 27].
@@
The logical gap (defined as the log-likelihood ratio difference between candidate corrections in distinct logical classes) has emerged as a Bayesian-motivated metric quantifying decoding confidence [8, 32, 62, 20]. Post-selection strategies based on the logical gap proved remarkably effective, enabling about 15-fold reductions in error rates with a relative overhead factor of <2<2 [8], error threshold improvements (up to 50% with full post-selection under code capacity noise) [62], and order-of-magnitude reductions in magic state preparation costs [33, 49]. The logical gap has become the standard approach for confidence-based post-selection in QEC.
Nevertheless, despite these advances, the logical gap method faces fundamental limitations that restrict broader applicability. First, computational overhead scales exponentially with the number $k$ of logical qubits (requiring comparative decoding over up to $4^{k}$ logical classes), which makes the method prohibitive for codes or circuits involving multiple logical qubits. Second, the method is not guaranteed to work well for general decoders beyond the optimal degenerate maximum likelihood decoder [41, 30] or the minimum-weight perfect matching (MWPM) decoder [28, 38, 37], as a correction within a fixed logical class may fail to represent the likelihood of the class.
@@ 2.1 Soft outputs from decoding
The minimum requirement of a decoder is to infer which logical correction should be applied to counteract errors, given the detector outcomes. Here, detectors indicate specific products of measurement outcomes deterministic in the absence of errors, which are normally constructed from two consecutive measurements of a stabilizer. A detector error model (DEM) [34] specifies the prior probabilities of independent fault locations (referred to as “error mechanisms”) and the detectors flipped by each of them. Decoders primarily aim to address the classification problem: given a DEM and detector outcomes, which logical class the current state belongs to. In practice, however, decoders typically produce additional information beyond this basic classification. These additional outputs are referred to as the soft outputs of the decoder.
@@
Here we focus on a specific application of soft outputs: quantifying decoding confidence. Specifically, we aim to estimate the reliability of the logical flip inferred by a decoder through analysis of its soft outputs. Given such a confidence metric, we construct a post-selection strategy that accepts a trial only when the metric value exceeds a specified cutoff threshold. We note that soft outputs have found broader applications beyond post-selection, including hierarchical code architectures where confidence information from inner codes guides the decoding of outer codes [32, 52, 63], construction of fast and accurate decoding framework by combining weak and strong decoders [67], and error mitigation for logical circuits [71, 27, 1]. However, this work focuses specifically on their application to post-selection.
@@ 2.1.1 Logical gap for quantifying decoding confidence
A prominent approach in the literature for quantifying decoding confidence involves comparative decoding and evaluation of the logical gap (also known as the complementary gap) [8, 32, 62]. Here, comparative decoding refers to a technique that performs decoding within each logical class and selects the minimum-weight correction among all classes. The logical gap is then defined as the difference between the two smallest log-likelihood weights of corrections in distinct logical classes.
To formalize the notion of a logical class, we fix an ordered set $\mathcal{L}=\{\bar{Q}_{1},\ldots,\bar{Q}_{l}\}$ of $l$ nontrivial independent logical Pauli errors. Each class is then labeled by a logical flip vector $\overrightarrow{\mathbf{\lambda}}\in\mathbb{Z}_{2}^{l}$, representing the logical effect of errors belonging to the class:
$\bar{P}\quantity(\overrightarrow{\mathbf{\lambda}})\coloneqq\prod_{j=1}^{l}\quantity(\bar{Q}_{j})^{\lambda_{j}}.$
Typically, $\mathcal{L}$ includes logical $X$ and $Z$ operators ($\bar{X}$, $\bar{Z}$) for each of $k$ logical qubits; however, some may act trivially on the circuit (e.g., a $\bar{Z}$ error just before a $\bar{Z}$ measurement) or may not be independent (e.g., $\bar{Z}_{1}$ and $\bar{Z}_{2}$ just before a $\bar{Z}_{1}\bar{Z}_{2}$ measurement), so $l\leq 2k$ in general. In our numerical analysis, we consider circuits involving only $\bar{Z}$ initialization/measurement, yielding $l=k$.
Comparative decoding aims to determine a correction conditioned on a chosen logical class $\overrightarrow{\mathbf{\lambda}}$. Operationally, this means we want to restrict decoding to errors whose net logical effect equals $\bar{P}(\overrightarrow{\mathbf{\lambda}})$. We implement this conditioning by augmenting the DEM with $l$ additional virtual detectors, one for each $\bar{Q}_{j}\in\mathcal{L}$, and then fixing their values to $\lambda_{j}$ during decoding. Intuitively, a virtual detector is an auxiliary syndrome bit that records whether a conjugate logical observable $\bar{R}_{j}$ would be flipped; by pinning this bit to $\lambda_{j}$, we force the decoder to output a correction in the desired logical class.
Formally, for each $\bar{Q}_{j}\in\mathcal{L}$ we choose a logical Pauli operator $\bar{R}_{j}$ anticommuting with $\bar{Q}_{j}$ and commuting with all other elements in $\mathcal{L}$. We then (conceptually) measure $\bar{R}_{j}$ immediately after the decoded circuit, obtaining an outcome from the product of certain physical measurement outcomes. The virtual detector associated with $\bar{Q}_{j}$ is defined to flip by every error mechanism that flips this product. [^1] If a logical measurement of $\bar{R}_{j}$ is already present in the circuit, we use it directly; otherwise, we temporarily append an ideal circuit fragment that measures $\bar{R}_{j}$ purely for the purpose of constructing the augmented DEM. Practically, this can be automated in Stim using OBSERVABLE_INCLUDE: (1) Wrap the circuit with the required logical $\bar{R}_{j}$ initialization/measurement layers when absent, (2) mark the corresponding measurement-outcome product as an observable via OBSERVABLE_INCLUDE, (3) generate the DEM, and (4) relabel the resulting observable as an additional detector so that it can be fixed to $\lambda_{j}$ during decoding. The temporary $\bar{R}_{j}$ measurement is an auxiliary construction and need not be included in actual comparative decoding runs. Decoding the augmented DEM with the virtual detectors fixed to $\overrightarrow{\mathbf{\lambda}}$ therefore yields a correction within the logical class $\overrightarrow{\mathbf{\lambda}}$. Note that virtual detectors generalize the notion of boundary detectors in the surface-code setting [32, 52].
After performing comparative decoding over all logical classes, the logical gap $\Delta$ is defined as the difference between the two smallest log-likelihood weights among the corrections $\{\tilde{E}_{\overrightarrow{\mathbf{\lambda}}}\}_{\overrightarrow{\mathbf{\lambda}}\in\mathbb{Z}_{2}^{l}}$ in different logical classes, i.e.,
$\Delta\coloneqq w_{(2)}-w_{(1)},$ (2)
@@
where $w_{(1)}\leq w_{(2)}\leq\cdots\leq w_{(2^{l})}$ are the sorted values of $w(\tilde{E}_{\overrightarrow{\mathbf{\lambda}}})]_{\overrightarrow{\mathbf{\lambda}}\in\mathbb{Z}_{2}^{l}}$. A large logical gap indicates that error patterns in alternative logical classes are significantly less likely than the predicted one, thereby implying high confidence in the decoding outcome.
Despite this general description, the logical gap method has been investigated almost exclusively for the minimum-weight perfect matching (MWPM) decoder. To understand why extending it to other decoders is fundamentally difficult, we first outline the core requirements that make the logical gap meaningful.
The logical gap quantifies how much more likely the predicted logical class is compared to alternatives. For this comparison to be valid, two conditions must hold: (i) The decoder must find a correction that faithfully represents the likelihood of each logical class, and (ii) the decoder must maintain its normal performance when constrained to decode within a specific logical class. The MWPM decoder satisfies both conditions: It can identify the minimum-weight (i.e., maximum-likelihood under independent noise) correction within any logical class, and its matching algorithm operates identically regardless of which logical class is specified. Other efficient decoders generally fail one or both of these requirements, as we elaborate below. [^2] We here focus on decoders with polynomial-time complexity, which are essential for practical QEC. In principle, computationally expensive decoders such as the degenerate quantum maximum likelihood decoder [25] can assess decoding confidence even more reliably than MWPM by evaluating the true likelihood of each logical class rather than approximating it via a single representative correction. However, their exponential computational cost in the general case renders them impractical for real-time decoding, despite recent progress on polynomial-time algorithms for restricted settings [16].
First, most efficient decoders besides MWPM do not guarantee finding the minimum-weight correction within a fixed logical class. When the decoder returns a suboptimal correction for a given class, the associated weight no longer accurately reflects that class’s likelihood. Consequently, comparing weights across classes becomes unreliable: A logical class might appear less likely simply because the decoder failed to find a good correction for it, not because errors in that class are genuinely improbable. This undermines the fundamental premise of the logical gap as a confidence measure.
Second, constraining a decoder to a specific logical class requires introducing “virtual” detectors that enforce the logical constraint (as described above). These virtual detectors can severely disrupt the internal heuristics of many decoders. For instance, the union-find (UF) decoder [23] grows clusters incrementally from violated detectors. A virtual detector, however, connects to $\Omega(d)$ error mechanisms (where $d$ is the code distance), causing any cluster touching it to grow abnormally fast and merge with distant clusters. This destroys the locality structure that the UF decoder exploits, leading to degraded decoding quality. The same issue affects generalized UF decoders for quantum low-density parity-check (QLDPC) codes [22], which rely on low-degree connectivity that high-degree virtual detectors violate.
@@
Due to these limitations, the logical gap method is generally not compatible beyond the MWPM decoder. However, the MWPM decoder is applicable only to a restricted class of QEC codes; specifically, those where each elementary error mechanism affects at most two detectors, such as repetition and surface codes [25]. A notable exception is the class of two-dimensional color codes [7], which are not directly compatible with MWPM but admit suboptimal decoders that apply MWPM to specific matchable subgraphs [24, 18, 5, 59, 46, 31, 48]. For these decoders, the logical gap method remains reasonably effective [49], particularly with the concatenated MWPM decoder [48].
@@
Last but not least, the logical gap method may be impractical due to its computational cost, which scales exponentially with the number $l$ of nontrivial independent logical Pauli errors (at most $2k$ for $k$ logical qubits), as it requires decoding across all $2^{l}$ logical classes. While manageable for memory experiments with a single logical qubit, this scaling renders the method difficult to use for multiple logical qubits with correlated noise. This limitation can be particularly fatal for QLDPC codes, which typically encode multiple logical qubits per code block; for instance, a single code block of the $[[144,12,12]]$ bivariate bicycle code [45, 11] has at most $4^{12}=$16777216$\$ logical classes.
@@ 2.2.1 Clustering-based decoders
To briefly introduce relevant terms first, the check matrix is a binary matrix indicating how error mechanisms (columns) and detectors (rows) are connected, which forms the biadjacency matrix of the bipartite Tanner graph with check nodes (corresponding to detectors) and fault nodes (corresponding to error mechanisms). A cluster is a set of error mechanisms that forms a connected region in the fault graph composed of only fault nodes, where two fault nodes are connected if the corresponding error mechanisms are commonly involved in at least one detector. A cluster is considered valid if it contains at least one solution consistent with all detectors that nontrivially involve any error mechanisms in the cluster. Formal definitions of these graph-theoretic structures are provided in Methods.
Figure 1: General workflow of a clustering-based decoder. (a) The tanner graph of the distance-9 surface code patch under bit-flip noise is shown with one possible example of detector outcome pattern. (b) A collection of disjoint valid clusters (which are guaranteed to have local solutions) are constructed according to a specific rule. The rule typically involves growing a cluster incrementally from each violated check node until it becomes valid, during which colliding clusters are merged. Note that while clusters are formally defined on the fault graph, we visualize them here on the Tanner graph since cluster growth rules naturally involve check node (detector) information. (c) Each cluster is decoded independently via a specific inner decoder. The local corrections are aggregated to form the final global correction. Note that the cluster construction and decoding processes shown here are for illustrative purposes only and do not correspond to any specific decoder.
As illustrated in Fig. 1, clustering-based decoders generally operate through three main steps (see Definition 8 in Methods for a formal definition):
  1. 1.
    Cluster construction: The decoder constructs a collection of disjoint valid clusters according to a specific rule, such that the union of detectors associated with these clusters covers all violated detectors. Typically, a cluster is grown incrementally from each violated check node until it becomes valid, during which colliding clusters are merged.
  2. 2.
    Local decoding: Each cluster is decoded independently by solving a reduced decoding problem restricted to the error mechanisms within the cluster and the detectors they affect.
  3. 3.
    Correction aggregation: The local corrections obtained from each cluster are combined to form the global correction.
The key differentiating factor among clustering-based decoders lies in how they organize the three components: (i) cluster growth, (ii) validity checking, and (iii) local decoding within each cluster.
The union-find (UF) decoder [23] is a representative example of this approach applicable to topological codes. In terms of the three components: (i) clusters are grown and fused incrementally from violated detectors using a union-find data structure, (ii) validity is checked by simple parity counting, and (iii) local decoding is performed via the peeling decoder which identifies a spanning tree for each cluster. The UF decoder was later generalized to quantum low-density parity-check (QLDPC) codes [22].
Another class of clustering-based decoders leverages belief propagation (BP) outcomes to guide cluster formation, prioritizing high-likelihood error mechanisms; examples include BP plus localized statistics decoding (BP+LSD) [39] and ambiguity clustering (AC) [69]. Specifically, the BP+LSD decoder first runs BP, accepting its solution if it converges, and otherwise applying LSD postprocessing. In LSD, (i) clusters are grown incrementally by including adjacent error mechanisms with highest BP likelihoods and (ii), (iii) a matrix factorization algorithm called on-the-fly elimination enables efficient validity check and local decoding [56, 53]. We provide a detailed description of BP+LSD in Methods, as it serves as the primary decoder throughout our numerical analysis.